With interest rates near an all-time low, a family decides to purchase their dream home. The house will cost $350,000. The family will pay 20% as a down payment, and finance the remaining balance with a 15-year fixed rate mortgage. The mortgage will call for monthly payments at a 4.50% APR.  How much interest is paid on the loan in its first two years?
Answer #1

EMI = P x r x (1+r) n/(1+r) n – 1

P = Loan amount = $ 350,000 x (1 - 0.2) = $ 350,000 x 0.8 = $ 280,000

r = Interest rate = 0.045/12 = 0.00375 p.m.

n = Number of periods = 15 x 12 = 180 periods

EMI = $ 280,000 x 0.00375 x (1+ 0.00375)180/ (1+ 0.00375) 180- 1

       = $ 1,050 x (1.00375)180/ (1.00375) 180- 1

       = $ 1,050 x 1.96155500814109 / 1.96155500814109 – 1

       = $ 2,059.63275854815 / 0.96155500814109

       = $ 2,141.9812086777 or $ 2,141.98

Loan amortization schedule for first two years:

Period

Beginning balance

(B)

Rate of Interest

(r)

EMI

Interest portion

(I = B x r)

Principal portion

(Pi = EMI - I)

Ending balance

(F = B - Pi)

1

$280,000

0.00375

$2,141.98

$1,050.00

$1,091.98

$278,908.02

2

278,908.02

0.00375

2,141.98

1,045.91

1,096.07

277,811.95

3

277,811.95

0.00375

2,141.98

1,041.79

1,100.19

276,711.76

4

276,711.76

0.00375

2,141.98

1,037.67

1,104.31

275,607.45

5

275,607.45

0.00375

2,141.98

1,033.53

1,108.45

274,499.00

6

274,499.00

0.00375

2,141.98

1,029.37

1,112.61

273,386.39

7

273,386.39

0.00375

2,141.98

1,025.20

1,116.78

272,269.61

8

272,269.61

0.00375

2,141.98

1,021.01

1,120.97

271,148.64

9

271,148.64

0.00375

2,141.98

1,016.81

1,125.17

270,023.47

10

270,023.47

0.00375

2,141.98

1,012.59

1,129.39

268,894.08

11

268,894.08

0.00375

2,141.98

1,008.35

1,133.63

267,760.45

12

267,760.45

0.00375

2,141.98

1,004.10

1,137.88

266,622.57

13

266,622.57

0.00375

2,141.98

999.83

1,142.15

265,480.43

14

265,480.43

0.00375

2,141.98

995.55

1,146.43

264,334.00

15

264,334.00

0.00375

2,141.98

991.25

1,150.73

263,183.27

16

263,183.27

0.00375

2,141.98

986.94

1,155.04

262,028.23

17

262,028.23

0.00375

2,141.98

982.61

1,159.37

260,868.85

18

260,868.85

0.00375

2,141.98

978.26

1,163.72

259,705.13

19

259,705.13

0.00375

2,141.98

973.89

1,168.09

258,537.05

20

258,537.05

0.00375

2,141.98

969.51

1,172.47

257,364.58

21

257,364.58

0.00375

2,141.98

965.12

1,176.86

256,187.72

22

256,187.72

0.00375

2,141.98

960.70

1,181.28

255,006.44

23

255,006.44

0.00375

2,141.98

956.27

1,185.71

253,820.74

24

253,820.74

0.00375

2,141.98

951.83

1,190.15

252,630.58

Total

$51,407.52

$24,038.10

$27,369.42

Total interest paid in first two years = $ 24,038.10